The lengths of the sides of a rectangle are consecutive odd integers. If the perimeter of the rectangle is 128 cm, what are the dimensions of the rectangle?
Answer by jorel1380(3719)

2n+2(n+2)=128
4n+4=128
4n=124
n=31
n+2=33
The rectangle is 33cm. by 31cm..